# DiffEyeSyn: Diffusion-based User-specific Eye Movement Synthesis

### v1

@@ 5.2. User Identity Recovery
Task Definition. We assume we have the identity-removed eye movements and the user embedding extracted from the ground truth 1,000 Hz eye movements. The objective is to recover the user identity for the identity-removed eye movements. It’s crucial to note that we achieved identity removal through downsampling, making this task equivalent to an extreme gaze data super-resolution task (from 20 Hz to 1,000 Hz) and an extreme gaze data imputation task with 98% missing values.
@@
Evaluation Metrics and Baselines. We assessed performance using cosine similarity between the EKYT embedding of the synthesised eye movements and that of the ground truth eye movements.
To the best of our knowledge, DiffEyeSyn is the first method for user-specific eye movement synthesis. Therefore, we compared it solely with its ablated version, which was trained without the proposed loss function for user identity guidance.
In addition, we established a human baseline to better explain the results. For each dataset, we computed within-user cosine similarity and cross-user cosine similarity. Specifically, within-subject cosine similarity measures how similar a sequence of eye movements is to other sequences from the same user. For each user’s sequence of eye movements, we calculated cosine similarities between its EKYT embedding and those of other sequences from the same user, then averaged these values across all users. Achieving an average cosine similarity close to the within-user cosine similarity suggests a method can simulate a user’s average gaze behaviour. An average cosine similarity higher than the within-user cosine similarity indicates that the method can simulate gaze behaviour similar to that of the given user embedding. For the Gazebase dataset, instead of only reporting the within-subject cosine similarity of the whole dataset, we also reported the within-subject cosine similarity of seven tasks.
@@
DiffEyeSyn w.o.$\mathcal{L}_{id}$ 0.404 ± 0.161 0.348 ± 0.173 0.352 ± 0.169 0.176 ± 0.211 0.269 ± 0.196 0.298 ± 0.193 0.407 ± 0.171 0.322 ± 0.182
GazeBase DiffEyeSyn 0.820 ± 0.063 0.800 ± 0.064 0.823 ± 0.079 0.705 ± 0.104 0.779 ± 0.081 0.791 ± 0.076 0.791 ± 0.081 0.787 ± 0.078
@@
DiffEyeSyn w.o.$\mathcal{L}_{id}$ 0.226 ± 0.126
JuDo1000 DiffEyeSyn - - - - - - - 0.695 ± 0.072
@@
Table 2. The cosine similarity between EKYT embeddings of the ground truth eye movement data and DiffEyeSyn synthesised data in user identity recovery task. For reference, the table shows the within-user and cross-user cosine similarities of both datasets. The best results among the models are bolded. Cosine similarity being 1 means that two embeddings are exactly the same, while 0 means two embeddings are orthogonal, i.e., not related.
@@
On the other hand, DiffEyeSyn without user identity guidance performs poorly in user identity recovery on both datasets. Specifically, on the GazeBase dataset, the cosine similarity consistently falls below the human within-user cosine similarity across all tasks by at least 0.06, with an overall cosine similarity 0.176 lower than the within-user cosine similarity. On the JuDo1000 dataset, it even performs worse than the cross-user cosine similarity. These findings highlight the failure of DiffEyeSyn without the proposed user identity guidance in injecting user-specific information into synthesised eye movements.
@@
In stark contrast, the full DiffEyeSyn surpasses the human within-user baseline on both datasets. For the GazeBase datasets, DiffEyeSyn outperforms the human within-user baseline by a considerable margin (at least 0.3) across all tasks. Despite the task of the JuDo1000 dataset being unseen during training, DiffEyeSyn achieves slightly better performance than the human within-user baseline (0.695 vs. 0.685), showcasing its strong generalisation ability. These results underscore DiffEyeSyn’s proficiency in reintroducing user-specific information into identity-removed eye movements.
@@
Qualitative Results. Due to the page limit, we present examples of synthesised eye movements in four different tasks (HSS, RAN, TEX, FXS) in Figure 3 and Figure 4. Figure 3 focuses on the full 5-second context and Figure 4 focuses more on micro eye movements with 100-millisecond duration. The tasks of TEX, along with others like VD1, VD2, and BLG, represent real-world situations, where choosing one exemplifies the general behavior adequately. Meanwhile, HSS, RAN, and FXS are chosen because they are challenging tasks that with certain types of eye events. More visualisations, including the visualisations of 5-second synthetic eye movements in all seven tasks, and visualisations of synthetic micro eye movements with a 100-millisecond time window, can be found in supplementary materials.
The synthetic eye movements generated by DiffEyeSyn for each task demonstrate high fidelity in replicating human eye behavior. Compared with the identity-removed inputs (left of each comparison), the synthesized outputs from DiffEyeSyn closely mirror the patterns of human eye movements, both in terms of velocity patterns and spatial trajectories. This similarity highlights the method’s capability to synthesise realistic user-specific eye movements.
@@
Figure 3. Qualitative comparison between the identity removed eye movements, DiffEyeSyn synthesised eye movements, and ground truth human eye movements in four tasks (HSS, RAN, TEX, FXS) within a 5-second time window. From left to right: identity removed eye movements, two DiffEyeSyn synthesised eye movements, human ground truth. For each sequence of eye movements, we visualise its velocities (above) and gaze direction (below).
Figure 4. Qualitative comparison between the identity removed eye movements, DiffEyeSyn synthesised eye movements, and ground truth human eye movements in four tasks (HSS, RAN, TEX, FXS) within a 100-millisecond time window. From left to right: identity removed eye movements, two DiffEyeSyn synthesised eye movements, human ground truth. For each sequence of eye movements, we visualise its velocities (above) and gaze direction (below).
@@ 5.3. User Identity Manipulation
Similar to the user identity recovery experiment, we downsampled the selected sequences to 20 Hz to evaluate our model performance on user-specific extreme gaze data super-resolution and gaze imputation.
Evaluation Metrics and Baselines. To evaluate model performance, we utilised cosine similarity between the EKYT embedding of the synthesised eye movements and that of the target eye movements. Similarly to the user identity recovery experiment detailed in Section 5.2, we established the ablated DiffEyeSyn, human within-user and cross-user cosine similarities as our baselines for comparison.
@@
Conversely, the full DiffEyeSyn surpasses the human within-user baselines across all tasks by a minimum of 0.05 and an average of 0.13. This indicates DiffEyeSyn’s capability to inject the user-specific information from the target user’s eye movement sequence into random eye movements effectively.
In the cross-dataset evaluation on the JuDo1000 dataset, DiffEyeSyn’s cosine similarity slightly trails the within-user baseline (0.605 vs. 0.685). However, it surpasses the cross-user similarity by 0.238, demonstrating DiffEyeSyn’s ability to synthesise user-specific eye movement behaviour that is distinguishable from others’ eye movement behaviour.
Dataset Method Task
HSS RAN TEX FXS VD1 VD2 BLG ALL
DiffEyeSyn w.o.$\mathcal{L}_{id}$ 0.235 ± 0.178 0.219 ± 0.179 0.209 ± 0.186 0.100 ± 0.169 0.164 ± 0.181 0.171 ± 0.183 0.214 ± 0.183 0.187 ± 0.179
GazeBase DiffEyeSyn 0.668 ± 0.101 0.664 ± 0.098 0.696 ± 0.106 0.454 ± 0.144 0.629 ± 0.129 0.644 ± 0.121 0.643 ± 0.110 0.628 ± 0.116
Human (within-user) - - - - - - - 0.685 ± 0.104
Human (cross-user) - - - - - - - 0.367 ± 0.157
DiffEyeSyn w.o.$\mathcal{L}_{id}$ 0.119 ± 0.126
JuDo1000 DiffEyeSyn - - - - - - - 0.605 ± 0.091
-- not applicable
Table 3. The cosine similarity between EKYT embeddings of the ground truth eye movement data and DiffEyeSyn synthesised data in user identity manipulation task. For reference, the table shows the within-user and cross-user cosine similarities of both datasets. The best results among the models are bolded. Cosine similarity being 1 means that two embeddings are exactly the same, while 0 means two embeddings are orthogonal, i.e., not related.
Qualitative Results. Figure 5 shows four examples of the user identity manipulation task. All the synthetic examples show more similar patterns to the target user’s eye movements in the velocity and spatial domain compared with the input eye movements. More examples can be found in supplementary materials.
Figure 5. Four examples of the user identity manipulation task. Left: the eye movements used to extract target user embedding. Middle: the eye movements from different users that DiffEyeSyn injects the target user information. Right: DiffEyeSyn synthesised eye movements. For each sequence of eye movements, we visualise its velocities (above) and gaze direction (below).
@@ 5.4. Comparison between Synthetic Velocity Distribution and Human Velocity Distribution
In eye movement synthesis research, prior studies often evaluate synthetic data quality by analysing statistics related to distinct eye events like fixations and saccades (Jiao et al., 2024; Prasse et al., 2023; Lan et al., 2022). Typically, this involves first detecting various eye events and then examining statistics such as fixation and saccade numbers, along with the velocity distributions of these events. Unlike these works, We opted not to directly compare statistics at the eye event level but instead focus on comparing velocity distributions between synthetic data and real human eye movements. This decision stemmed from our user identity recovery experiment (see Section 5.2), where our emphasis was on reintroducing user-specific information into input eye movements. Given this context, the absolute count of eye events does not vary a lot from actual human eye movements. On the other hand, our method synthesises eye movement velocities and subsequently translates these velocities into gaze directions. Notably, common eye event detection algorithms like I-VT (Salvucci and Goldberg, 2000) and I-DT (Salvucci and Goldberg, 2000) operate based on velocity thresholds or dispersion criteria. Thus, for our synthetic data, identifying eye events is equivalent to detecting them based on eye movement velocities and predefined thresholds. This rationale supports our decision to forego eye event-level statistics in favour of direct velocity distribution comparisons between synthetic and human eye movements.
Evaluation Metrics and Baselines. Following prior works (Prasse et al., 2023; Manning and Schutze, 1999), we used the Jensen-Shannon divergence to measure the similarity between two distributions. The Jensen–Shannon divergence ($JS$) is a symmetrised and smoothed version of the Kullback–Leibler divergence ($KL$). Given discrete probability distributions $P$ and $Q$, the $JS$ is defined as:
@@
We established a strong human baseline by calculating the $JS$ between the human ground truth of the GazeBase test set and the JuDo1000 dataset.
Exprimental Details. Given the large scale of the two datasets we used, directly comparing distributions from all samples in the dataset becomes infeasible. To address this, we adopted a sampling approach by selecting 1,000,000 velocity samples from each data distribution for distribution comparison. This process was repeated ten times to mitigate sampling bias, and we recorded the average $JS$ value (marked as ALL in 4).
Additionally, we applied the I-VT (Salvucci and Goldberg, 2000) algorithm for fixation and saccade detection. Eye movements with velocities below $100^{\circ}/s$ were classified as fixations, while those exceeding $300^{\circ}/s$ were categorized as saccades. Consistent with our approach for calculating the average JS divergence between the velocity distributions of all synthetic data and human ground truth, we also computed and reported the JS divergence between the velocity distributions of synthetic fixation samples and human fixations, as well as between synthetic saccade samples and human saccades.
Quantitative Results. The results are shown in Table 4. The velocity distributions of the synthetic eye movements across two different tasks and two different datasets achieve lower $JS$ values than the human baseline (comparing the velocity distributions between the JuDo1000 dataset and the GazeBase dataset). This indicates that DiffEyeSyn generated human-level realistic eye movements from a perspective of velocity distribution. Specifically, the velocity distribution of DiffEyeSyn generated eye movements is closer to the velocity distribution of target users than other humans.
Method Distribution 1 Distribution 2 Jensen-Shannon divergence(×$\times 10^{-2}$) $\downarrow$
@@
DiffEyeSyn Synthetic GazeBase (Recovery) GazeBase 1.35 0.35 0.35
Synthetic GazeBase (Manipulation) GazeBase 1.03 0.27 0.40
Synthetic JuDo1000 (Recovery) JuDo1000 2.31 0.08 1.17
Synthetic JuDo1000 (Manipulation) JuDo1000 2.75 0.35 1.61
Human JuDo1000 GazeBase 6.96 5.80 3.89
GazeBase GazeBase 0 0 0
JuDo1000 JuDo1000 0 0 0
Table 4. Jensen-Shannon divergence between human eye movement data and DiffEyeSyn synthetic eye movement data from experiments mentioned in Section 5.2 and 5.3. Bold values indicate the results of the method are lower than one of the human baselines.

### v2

@@ 5.2. User Identity Recovery
Task Definition. We assume we have the identity-removed eye movements and the user embedding extracted from the ground truth 1,000 Hz eye movements. The objective is to recover the user identity for the identity-removed eye movements.
@@
Evaluation Metrics and Baselines. We assessed performance using cosine similarity between the EKYT embedding of the synthesised eye movements and that of the ground truth eye movements. We reported the mean cosine-similarity and the standard deviation.
To the best of our knowledge, DiffEyeSyn is the first method for user-specific eye movement synthesis. We compared it with its ablated version, which was trained without the proposed loss function for user identity guidance.
We also compared DiffEyeSyn with a classical signal processing baseline. Since we removed user identity by downsampling eye movements to 20 Hz, this is equivalent to removing the high-frequency components. For the baseline, we applied a Butterworth high-pass filter to extract frequency components above 20 Hz from the ground truth, then added these back to the identity-removed eye movements.
In addition, we benchmarked against SP-EyeGAN (Prasse et al., 2023, 2024), an existing eye movement synthesis method. It requires predefined fixation locations, mean, and standard deviation for fixation and saccade durations as input for eye movement synthesis. We trained one SP-EyeGAN model for each task in the GazeBase dataset using their official code. For each eye movement sequence in the GazeBase, we synthesised one eye movement sequence using the corresponding SP-EyeGAN model. For cross-dataset evaluation on the JuDo1000 dataset, we employed the SP-EyeGAN model trained on the RAN task.
Moreover, we established a human baseline to better explain the results. For each dataset, we computed within-user cosine similarity and cross-user cosine similarity. Specifically, within-subject cosine similarity measures how similar a sequence of eye movements is to other sequences from the same user. For each user’s sequence of eye movements, we calculated cosine similarities between its EKYT embedding and those of other sequences from the same user, then averaged these values across all users. Achieving an average cosine similarity close to the within-user cosine similarity suggests a method can simulate a user’s average gaze behaviour. An average cosine similarity higher than the within-user cosine similarity indicates that the method can simulate gaze behaviour similar to the target eye movement sequence. For the Gazebase dataset, we reported the within-subject cosine similarity of seven tasks and the whole dataset.
@@
SP-EyeGAN (Prasse et al., 2023, 2024) 0.114 ± 0.136 0.134 ± 0.137 0.157 ± 0.142 0.106 ± 0.150 0.128 ± 0.145 0.121 ± 0.144 0.135 ± 0.133 0.129 ± 0.141
High-pass Filter (R) 0.390 ± 0.170 0.407 ± 0.178 0.320 ± 0.155 0.477 ± 0.241 0.362 ± 0.178 0.362 ± 0.174 0.324 ± 0.178 0.375 ± 0.180
DiffEyeSyn w.o.$\mathcal{L}_{id}$ (R) 0.404 ± 0.161 0.348 ± 0.173 0.352 ± 0.169 0.176 ± 0.211 0.269 ± 0.196 0.298 ± 0.193 0.407 ± 0.171 0.322 ± 0.182
GazeBase DiffEyeSyn (R) 0.820 ± 0.063 0.800 ± 0.064 0.823 ± 0.079 0.705 ± 0.104 0.779 ± 0.081 0.791 ± 0.076 0.791 ± 0.081 0.787 ± 0.078
High-pass Filter (M) 0.181 ± 0.148 0.178 ± 0.151 0.178 ± 0.157 0.223 ± 0.143 0.190 ± 0.152 0.194 ± 0.150 0.153 ± 0.145 0.182 ± 0.151
DiffEyeSyn w.o.$\mathcal{L}_{id}$ (M) 0.235 ± 0.178 0.219 ± 0.179 0.209 ± 0.186 0.100 ± 0.169 0.164 ± 0.181 0.171 ± 0.183 0.214 ± 0.183 0.187 ± 0.179
DiffEyeSyn (M) 0.668 ± 0.101 0.664 ± 0.098 0.696 ± 0.106 0.454 ± 0.144 0.629 ± 0.129 0.644 ± 0.121 0.643 ± 0.110 0.628 ± 0.116
@@
SP-EyeGAN (Prasse et al., 2023, 2024) - - - - - - - 0.141 ± 0.136
High-pass Filter (R) - - - - - - - 0.534 ± 0.162
DiffEyeSyn w.o.$\mathcal{L}_{id}$ (R) - - - - - - - 0.226 ± 0.126
JuDo1000 DiffEyeSyn (R) - - - - - - - 0.695 ± 0.072
High-pass Filter (M) - - - - - - - 0.424 ± 0.139
DiffEyeSyn w.o.$\mathcal{L}_{id}$ (M) - - - - - - - 0.119 ± 0.126
DiffEyeSyn (M) - - - - - - - 0.605 ± 0.091
@@
Table 2. The cosine similarity between EKYT embeddings of the ground truth eye movement data and synthetic data of different methods in user identity recovery (R) and user identity manipulation (M). The results that are better than the human within-user baseline the models are bolded. Cosine similarity being 1 means that two embeddings are exactly the same, while 0 means two embeddings are orthogonal, i.e., not related.
@@
In comparison, SP-EyeGAN achieves performance close to the cross-user human baseline on the GazeBase dataset and performs even worse than this baseline on the JuDo1000 dataset, suggesting that it struggles to capture user-specific characteristics in eye movements.
@@
On the other hand, although significantly outperforming SP-EyeGAN, DiffEyeSyn without user identity guidance performs poorly in user identity recovery on both datasets. Specifically, on the GazeBase dataset, the cosine similarity consistently falls below the human within-user cosine similarity across all tasks by at least 0.06, with an overall cosine similarity 0.176 lower than the within-user cosine similarity. On the JuDo1000 dataset, it even performs worse than the cross-user cosine similarity. These findings highlight the failure of DiffEyeSyn without the proposed user identity guidance in injecting user-specific information into synthesised eye movements.
The high-pass filter baseline slightly outperforms the ablated DiffEyeSyn on both datasets but still underperforms relative to the within-user baseline. This finding suggests that user-specific information resides not only in the high-frequency components (above 20 Hz) but also in the low-frequency components of eye movement sequences.
@@
In stark contrast, the full DiffEyeSyn surpasses the human within-user baseline on both datasets. For the GazeBase datasets, DiffEyeSyn outperforms the human within-user baseline by a considerable margin (at least 0.3) across all tasks. Despite the task of the JuDo1000 dataset being unseen during training, DiffEyeSyn achieves slightly better performance than the human within-user baseline (0.695 vs. 0.685), showcasing its strong generalisation ability. These results underscore DiffEyeSyn’s proficiency in reintroducing user-specific subtle eye movements into identity-removed eye movements.
@@
Qualitative Results. Due to the page limit, we present examples of synthesised eye movements in four different tasks (HSS, RAN, TEX, FXS) in Figure 4. The tasks of TEX, along with others like VD1, VD2, and BLG, represent real-world situations, where choosing one exemplifies the general behavior adequately. Meanwhile, HSS, RAN, and FXS are chosen because they are challenging tasks that with certain types of eye events. More visualisations, including the visualisations of 5-second synthetic eye movements in all seven tasks can be found in the supplementary materials.
Compared with the identity-removed inputs and the synthetic eye movements generated by the high-pass filter, DiffEyeSyn demonstrates high fidelity in replicating human eye movement velocities across all tasks. In addition, the high-pass filter approach suffers from numerical instability, which can result in invalid output values (see the last two rows in Figure 4). In stark contrast, DiffEyeSyn is robust in synthesising user-specific eye movements and is capable of injecting user-specific information into any given sequence of eye movements.
@@
Figure 4. Qualitative comparison between the identity removed eye movements, High-pass filter synthetic eye movements, DiffEyeSyn synthesised eye movements, and ground truth human eye movements in four tasks (HSS, RAN, TEX, FXS) within a 5-second time window. For each sequence of eye movements, we visualise its velocities (above) and gaze direction (below). A figure is empty means the method fails to produce valid results.
@@ 5.3. User Identity Manipulation
Similar to the user identity recovery experiment, we downsampled the selected sequences to 20 Hz to remove the user-specific information.
Evaluation Metrics and Baselines. To evaluate model performance, we utilised cosine similarity between the EKYT embedding of the synthesised eye movements and that of the target eye movements. Similarly to the user identity recovery experiment detailed in Section 5.2, we established the ablated DiffEyeSyn, SP-EyeGAN, high-pass filter, human within-user and cross-user cosine similarities as our baselines for comparison. For the high-pass filter baseline in this experiment, we extracted frequency components above 20 Hz from the target user’s eye movements and injected them into the downsampled base sequences.
@@
In the Gazebase dataset, the performance of the high-pass filter dropped a lot compared with the performance in the user-identity recovery task; it was only slightly higher than the cross-user baseline. Conversely, the full DiffEyeSyn surpasses the human within-user baselines across all tasks by a minimum of 0.05 and an average of 0.13. This indicates DiffEyeSyn’s capability to inject the user-specific information from the target user’s eye movement sequence into random eye movements effectively.
In the cross-dataset evaluation using the JuDo1000 dataset, DiffEyeSyn achieves a cosine similarity of 0.605, trailing slightly behind the within-user baseline of 0.685. However, it surpasses the cross-user baseline by a margin of 0.238, demonstrating its ability to synthesise eye movement patterns that are meaningfully aligned with the target user’s identity. The high-pass filter baseline achieved a cosine similarity of just 0.424, which is 0.181 lower than DiffEyeSyn, further highlighting the advantage of our proposed method.
Qualitative Results. Figure 5 presents four examples from the user identity manipulation task. As with the results in Figure 4, the high-pass filter consistently fails to produce valid outputs in some cases. Furthermore, its generated eye movement velocities appear considerably noisier than those of the target user, leading to unrealistic jitter and instability in the synthesised gaze directions. In contrast, DiffEyeSyn demonstrates greater robustness and produces synthetic eye movements that are visually more realistic, both in terms of velocity profiles and spatial patterns. Additional examples are provided in the supplementary materials.
To further understand how the base and target eye movements contribute to the synthesis process, we present four additional examples in Figure 6, where the same base sequence is used with different target user embeddings. All DiffEyeSyn synthesised results are spatially similar to the base sequences, but show subtle, meaningful differences. This suggests that the base sequence constrains the overall movement range and high-level eye movement structure (e.g., the number, type, and order of events), while the injected user embedding refines finer details, such as producing varying saccade amplitudes at the same saccade locations.
Figure 5. Four examples of the user identity manipulation task. Left: the eye movements used to extract the target user embedding. Middle 1: the eye movements from different users that DiffEyeSyn injects the target user information. Middle 2: High-pass filter synthesised eye movements. Right: DiffEyeSyn synthesised eye movements. For each sequence of eye movements, we visualise its velocities (above) and gaze direction (below). A figure is empty means the method fails to produce valid results.
Figure 6. Four examples of the user identity manipulation task with the same base eye movement sequence. Left: the eye movements used to extract the target user embedding. Middle: the eye movements that DiffEyeSyn injects the target user information. Right: DiffEyeSyn synthesised eye movements. For each sequence of eye movements, we visualise its velocities (above) and gaze direction (below).
@@ 5.4. Comparison between Synthetic Velocity Distribution and Human Velocity Distribution
In eye movement synthesis research, prior studies often evaluate synthetic data quality by analysing statistics related to distinct eye events like fixations and saccades (Jiao et al., 2024; Prasse et al., 2023; Lan et al., 2022). Typically, this involves first detecting various eye events and then examining statistics such as fixation and saccade numbers, along with the velocity distributions of these events. Unlike these works, We opted not to directly compare statistics at the eye event level but instead focus on comparing velocity distributions between synthetic data and real human eye movements. This decision stemmed from our user identity recovery and manipulation experiments, where our emphasis was on reintroducing user-specific information into input eye movements. Given this context, the absolute count of eye events does not vary a lot from actual human eye movements. On the other hand, our method synthesises eye movement velocities and subsequently translates these velocities into gaze directions. Notably, common eye event detection algorithms like I-VT (Salvucci and Goldberg, 2000) and I-DT (Salvucci and Goldberg, 2000) operate based on velocity thresholds or dispersion criteria. Thus, for our synthetic data, identifying eye events is equivalent to detecting them based on eye movement velocities and predefined thresholds. This rationale supports our decision to forego eye event-level statistics in favour of direct velocity distribution comparisons between synthetic and human eye movements.
Evaluation Metrics and Baselines. Following prior works (Prasse et al., 2023, 2024; Manning and Schutze, 1999), we used the Jensen-Shannon divergence to measure the similarity between two distributions. The Jensen–Shannon divergence ($JS$) is a symmetrised and smoothed version of the Kullback–Leibler divergence ($KL$). Given discrete probability distributions $P$ and $Q$, the $JS$ is defined as:
@@
We evaluated DiffEyeSyn against SP-EyeGAN and the high-pass filter baseline using the synthesised data described in Section 5.
Exprimental Details. Given the large scale of the two datasets we used, directly comparing distributions from all samples in the dataset becomes infeasible. To address this, we adopted a sampling approach by selecting 1,000,000 velocity samples from each data distribution for distribution comparison. This process was repeated ten times to mitigate sampling bias, and we recorded the average $JS$ value (marked as ALL in Table 3).
Additionally, we employed the I-VT algorithm (Salvucci and Goldberg, 2000) to identify fixations and saccades. Specifically, eye movements with velocities below $100^{\circ}/\text{s}$ were classified as fixations, while those exceeding $300^{\circ}/\text{s}$ were categorised as saccades. In line with our approach for computing the average JS divergence between the velocity distributions of all synthetic samples and the human ground truth, we also calculated and reported the JS divergence between the velocity distributions of synthetic fixations and real fixations, as well as between synthetic and human saccades.
This additional evaluation is motivated by the fact that the majority of human eye movements fall within the fixation category, which generally involves slower velocities. Consequently, a method might achieve favourable JS scores across all samples by simply modelling low-velocity fixations well, while failing to capture the more dynamic saccadic behaviour. By examining fixations and saccades separately, we gain a clearer picture of the method’s ability to model the full spectrum of eye movement velocities.
To compute these distributions, we used the histogram function from NumPy. As velocities were clipped at a maximum of $1000^{\circ}/\text{s}$, we selected bins = 500 for all samples, bins = 350 for saccades, and bins = 50 for fixations. This ensured a consistent bin width of 2 across all histograms, allowing for direct comparison between JS divergence scores.
Quantitative Results. The results are summarised in Table 3. Across both tasks and datasets, the velocity distributions of synthetic eye movements generated by DiffEyeSyn consistently yield the lowest $JS$ divergence scores when compared with those produced by SP-EyeGAN and the high-pass filter baseline. This indicates that DiffEyeSyn generates synthetic data that most closely approximates the velocity distribution of real human eye movements.
Notably, DiffEyeSyn retains this performance advantage not only when considering the full set of samples but also when fixations and saccades are evaluated separately. This suggests that DiffEyeSyn is capable of faithfully modelling both slow and rapid eye movement behaviours—an essential quality for synthesising realistic and nuanced gaze data. In contrast, the high-pass filter exhibits markedly higher divergence across all categories, reflecting its limited capacity to replicate human-like eye movement dynamics, which is consistent with the qualitative observations shown in Figure 5. While SP-EyeGAN achieves comparable performance to DiffEyeSyn on fixation data, it performs noticeably worse in modelling saccadic movements, further highlighting the strengths of our approach in capturing fast eye movement behaviour.
Dataset Synthetic Eye Movements JS divergence (×$\times 10^{-2}$) $\downarrow$
@@
GazeBase SP-EyeGAN (Prasse et al., 2023, 2024) 1.28 7.98 1.59
Hiph-pass Filter (R) 15.68 4.68 15.43
DiffEyeSyn (R) 1.10 0.35 1.05
Hiph-pass Filter (M) 18.44 3.94 18.17
DiffEyeSyn (M) 1.01 0.27 1.00
JuDo1000 SP-EyeGAN (Prasse et al., 2023, 2024) 5.79 2.87 4.63
Hiph-pass Filter (R) 6.66 4.35 6.69
DiffEyeSyn (R) 2.26 0.10 2.17
Hiph-pass Filter (M) 7.25 8.81 7.43
DiffEyeSyn (M) 2.69 0.36 2.59
Table 3. Jensen-Shannon divergence between velocity distribution of ground truth human eye movements and velocity distribution of synthetic eye movements of different methods from experiments mentioned in Section 5.2 and 5.3. The best results are bolded and the second best are underlined. R: user identity recovery, M: user identity manipulation.
5.5. Augmenting Gaze-based User Identification
To further evaluate the effectiveness of DiffEyeSyn in capturing user-specific information, we conducted an additional gaze-based user identification experiment involving 59 users from the GazeBase test set. In this experiment, we augmented the training data by incorporating synthetic eye movements generated by DiffEyeSyn. If the synthetic data fails to preserve user-specific characteristics, the identification model trained on the augmented dataset would be expected to underperform compared to a model trained solely on the original data, as previously reported in (Jiao et al., 2023).
Experimental Details. For each user in the GazeBase test set, we randomly selected 50% eye movement trajectories for training, and the rest for testing. Additionally, we created a DiffEyeSyn-augmented training set by synthesising seven new samples for each original training instance, as described in Section 5.3.
We employed the EKYT model for user identification, extending it with a final linear classification layer. The models were trained separately on the original human data and the DiffEyeSyn-augmented dataset. Both models were trained for 100 epochs using the Adam optimiser, with a learning rate of 0.001 and a batch size of 128. To minimise training variability, we ran each model five times with different random seeds. In line with prior work (Jiao et al., 2023), we reported the average classification accuracy and equal error rate (EER) as evaluation metrics.
Training data Acc. (%) $\uparrow$ EER (%) $\downarrow$
Human 57.77 ± 2.98 1.72 ± 0.19
Augmented 59.63 ± 1.72 1.70 ± 0.20
Table 4. User identification results of the same model trained on the Gazebase (Human) dataset and the DiffEyeSyn augmented dataset. The best results are bolded. Acc: classification accuracy.
Table 4 shows that the model trained on the dataset augmented with DiffEyeSyn achieves superior performance across both evaluation metrics. This finding suggests that DiffEyeSyn effectively captures user-specific features during synthesis, thereby enhancing the utility of existing gaze datasets for downstream tasks.